Q1 Improper Integrals 10 Points Evaluate the following integrals and determine if they converge. If they converge, find the vshow works please
Answer #1

Gian ļ at der gte? Subi Ա - du ze dy Incu) n 다 du น U-27e7 - In cate%) = ln (ate) = ln(at ea In (3) In(a) =) (3) -In(a) = Con